# Probing Reheating Phase via Non-Helical Magnetogenesis and Secondary Gravitational Waves

### v1

@@ III.3.2 Spectral Analysis of Secondary Gravitational Waves (SGWs)
In scenarios with infinite electrical conductivity during reheating, the electric fields vanish, implying that any contributions from electric fields originate solely from inflationary production. For the case $n<-1/2$, where electric fields dominate, the spectrum follows Eq. (58) for both $w_{re}<1/3$ and $w_{re}>1/3$ reheating scenarios.
Conversely, for $n>1/2$, magnetic fields dominate, and the resulting spectrum follows the form given in Eq. (58) for $w_{re}<1/3$, or Eq. (III.3.2) for $w_{re}>1/3$.
Figure 2: In this figure, we plot $\Omega_{\mathrm{gw}}h^{2}$ versus $f$ (in Hz) for $w_{re}=0$ for both scenarios, $n>1/2$ and $n<-1/2$. In the left panel, we consider different values of the magnetic or electric spectral index while keeping the reheating temperature fixed at $T_{\rm re}=10^{3}$ GeV. In the right panel, we show the dependence of the GW spectra on the reheating temperature for a fixed value of the magnetic (or electric) spectral index.
Figure 3: In the above figure, we plot $\Omega_{\mathrm{gw}}h^{2}$ as a function of frequency $f$ (in Hz) for magnetogenesis scenarios with $n>1/2$, where the magnetic field strength is greater than that of the electric field. In the left panel, we show how the spectrum depends on the magnetic spectral index for a fixed reheating scenario with $w_{re}=3/5$ and $T_{\rm re}=10^{6}$ GeV. In the right panel, we illustrate the dependence of the reheating temperature on the GW spectrum for a fixed value of $n_{\mathrm{B}}=0.2$ and $w_{re}=3/5$.
Figure 4: In the figures above, we have plotted $\Omega_{\mathrm{gw}}h^{2}$ as a function of frequency $f$ (in Hz) for the coupling $n<-1/2$, assuming zero electrical conductivity during the reheating era. In the left panel, we show the dependence of the coupling parameter $n~{}(n_{\mathrm{E}})$ for a fixed reheating scenario with $w_{re}=3/5$ and $T_{\rm re}=10^{6}$ GeV. In the right panel, we display the dependence of the reheating temperature on the GW spectrum for a fixed equation of state $w_{re}=3/5$ and a fixed magnetogenesis coupling parameter $n=1.9~{}(n_{\mathrm{E}}=0.2)$.
@@ IV Results and Discussion:
In Fig. (2), we plot $\Omega_{\mathrm{gw}}h^{2}$ as a function of present-day observable frequency $f$ (Hz) for a reheating scenario with $w_{re}=0$, where the universe behaves as a matter-like fluid. The left panel illustrates the dependency on the initial electromagnetic profile for couplings $n>1/2$ and $n<-1/2$, while the right panel shows dependency on the duration of reheating. For $n>1/2$, where magnetic fields dominate during inflation, the entire SGW spectrum originates from these fields. In contrast, for $n<-1/2$, the electric field dominates, making the GW spectral index dependent on the electric spectral index $n_{\mathrm{E}}$. Both fields show similar spectral behavior under a sign change in $n$, resulting in similar GW spectra when $w_{re}<1/3$, where inflationary GW production dominates. On large scales, the spectrum is scale-invariant, while sub-horizon modes re-entering before reheating end show a red-tilted spectrum for $f>f_{\rm re}$, scaling as $f^{2n_{\mathrm{B}}-|n_{w}|}$ for $w_{re}<1/3$. In the plot we consider $w_{re}=0$ for which $n_{w}=2(1-3w_{re})/(1+3w_{re})=2$. Increasing the reheating temperature effectively shortens the duration of the reheating phase, thereby shifting the peak frequency of the gravitational wave spectrum toward higher values. In both figures, the large-wavelength (low-frequency) modes exhibit a scale-invariant gravitational wave spectrum, i.e., $\Omega_{\mathrm{gw}}h^{2}(f\ll f_{\rm re})\propto f^{0}$. This behavior arises from the contribution of primordial gravitational waves (PGWs). Since we consider blue-tilted electromagnetic (EM) fields, most of the magnetic energy is stored in high-frequency modes. At low frequencies, however, the dominant contribution comes from PGWs, which are generated from vacuum fluctuations during inflation and possess a scale-invariant spectrum (see Eq. (57)).
In Fig.(3), we plot $\Omega_{\mathrm{gw}}h^{2}$ for $w_{re}>1/3$, with coupling $n>1/2$ where the magnetic field is dominant. Larger values of $n_{\mathrm{B}}$ yield spectra that can exceed upper bounds set by $\Delta N_{\rm eff}$ constraintsClarke et al. (2020a). Lower values of $n_{\mathrm{B}}$, however, are consistent with $\Delta N_{\rm eff}$ and tensor-to-scalar ratio bounds and introduce additional spectral breaks. In the range $f_{\nu}<f<f_{\rm re}$, SGWs from the radiation-dominated era exceed those produced during reheating. At high frequencies, $f\gg f_{\rm re}$, the spectrum follows $f^{2n_{\mathrm{B}}+|n_{w}|}$ as inflationary SGW production dominates. The left panel shows the dependence on the magnetic spectral index for $w_{re}=3/5$ with $T_{\rm re}=10^{6}$ GeV, while the right panel examines reheating duration dependency for fixed $w_{re}=3/5$ and $n_{\mathrm{B}}=0.2$. As reheating temperature decreases, the GW amplitude is enhanced by $(k_{\mathrm{re}}/k_{\rm e})^{|n_{w}|}$, as the background energy density dilutes faster than the GW density. At very high frequencies, i.e., $f\gg f_{\rm re}$, the gravitational wave spectrum follows a simple scaling, $\Omega_{\mathrm{gw}}h^{2}\propto f^{0.57}$, for all parameter sets, differing only in amplitude. This behavior arises from the contribution of primordial gravitational waves (PGWs) at these scales. As is well known, for $w_{re}>1/3$, the PGW spectrum is blue-tilted, scaling as $\Omega_{\mathrm{gw}}^{\mathrm{pri}}(f)\propto f^{-n_{w}}$ (see Eq. (57)). Since we have fixed $w_{re}=3/5$, the spectral index at very high frequencies is determined by $n_{w}$. The overall amplitude, however, varies depending on the reheating temperature. A lower reheating temperature implies that sub-horizon modes spend a longer time during the reheating phase, leading to greater enhancement of their amplitudes compared to scenarios with higher reheating temperatures.
In Fig. 4, we present the gravitational wave (GW) spectrum $\Omega_{\mathrm{gw}}h^{2}$ for $n<-1/2$, where electric fields dominate during inflation. Assuming zero electrical conductivity during reheating, electric fields remain unsuppressed and significantly contribute to GW production in the intermediate frequency range. In this scenario, there are two distinct phases during which electric fields can source gravitational waves: (i) the inflationary epoch and (ii) the reheating era, enabled by the assumption of vanishing conductivity. After reheating ends, the universe enters the standard radiation-dominated phase, during which the plasma conductivity becomes large. Consequently, electric fields—especially those on sub-horizon scales—are rapidly dissipated due to the prompt response of the charged particles, and thus contribute negligibly to the GW background in this era, even when considering reheating scenarios with $w_{re}>1/3$. The shape of the spectrum in this case is characteristically distinct. At high frequencies, the dominant contribution arises from primordial gravitational waves (PGWs), and the spectrum follows $\Omega_{\mathrm{gw}}h^{2}\propto f^{-n_{w}}$. At low frequencies ($f<f_{\rm re}$), the spectrum exhibits a blue tilt, $\Omega_{\mathrm{gw}}h^{2}\propto f^{2n_{\mathrm{E}}}$, which is only possible if secondary gravitational waves (SGWs), sourced by the electric fields, dominate over the PGWs. This condition depends sensitively on the inflationary coupling parameter $n$, as well as the reheating parameters $w_{re}$ and $T_{\rm re}$. Furthermore, for certain parameter choices, we observe that the GW spectrum remains scale-invariant at CMB scales, indicating PGW dominance at those wavelengths. Additionally, decreasing the reheating temperature for fixed $w_{re}$ and $n_{\mathrm{E}}$ (see the right panel of Fig. 4) leads to an overall enhancement in GW amplitude, due to the prolonged duration of reheating. This also shifts the spectral break near $f\sim f_{\rm re}$ toward lower frequencies.
@@
Using the limit $r_{0.05}\leq 0.036$, we find that the coupling parameter $n$ for magnetogenesis should satisfy $-2.1\leq n\leq 2.1$. This bound holds universally, even for instantaneous reheating. For prolonged reheating scenarios with $w_{re}>1/3$, these constraints become more robust, limiting both the reheating dynamics and the magnetogenesis model. Thus, utilizing the upper bound $r_{0.05}\leq 0.036$ at the pivot scale, we derive the following constraint,
$1\geq\frac{182~{}rA_{s}\mathcal{A}_{1}}{(3-2n_{\mathrm{B}})}\left(\frac{k_{\rm e}}{k_{\mathrm{re}}}\right)^{2|n_{w}|}\left(\frac{k_{*}}{k_{\rm e}}\right)^{2n_{\mathrm{B}}}\mathcal{C}_{{}_{\rm B/E}}^{2}(n_{{}_{\mathrm{B}/\mathrm{E}}}).$ (61)
Using this constraint, we set bounds on the reheating temperature for specific values of the equation of state (EoS) across five discrete values of the magnetic spectral index $n_{{}_{\mathrm{B}/\mathrm{E}}}=0.01,0.1,0.2,0.3,\text{ and }0.4$, as shown in Table (2).
$w_{re}$ $n_{\mathrm{B}}=0.01$ $n_{\mathrm{B}}=0.1$ $n_{\mathrm{B}}=0.2$ $n_{\mathrm{B}}=0.3$ $n_{\mathrm{B}}=0.4$
@@
0.4 $1.77\times 10^{9}$- -- -- --
0.5 $5.65\times 10^{12}$ $1.28\times 10^{7}$ $2.89$- --
0.6 $8.2\times 10^{13}$ $1.46\times 10^{10}$ $5.41\times 10^{5}$- --
@@
Table 2: In the table above, we’ve listed the minimum reheating temperature for various equations of states (denoted as $w_{re}$), considering five different values of the magnetic spectral index (denoted as $n_{\mathrm{B}}$).
In our analysis, we treat the reheating equation-of-state parameter $w_{re}$ as a free parameter, allowing it to take arbitrary values. However, in specific inflationary models—for example, those with power-law potentials of the form $V(\phi)\propto\phi^{2n}$, $w_{re}$ can be uniquely determined based on the dynamics of the scalar field. In this work, we adopt a generalized approach by leaving $w_{re}$ unconstrained. As an illustration, to avoid the overproduction of tensor perturbations at CMB scales, consider the case where the magnetic spectral index is $n_{\mathrm{B}}=0.01$ (nearly scale-invariant). For a reheating equation-of-state $w_{re}=0.4$, the lower bound on the reheating temperature is found to be $T_{\rm re}^{\rm min}\simeq 1.77\times 10^{9}\,\mathrm{GeV}$. If the reheating temperature falls below this value, electromagnetic fields generated via this mechanism can produce excessive tensor fluctuations at large scales, violating current observational constraints on the tensor-to-scalar ratio. We summarize the lower bounds on reheating temperatures for various parameter combinations in Table (2).
@@
During the epoch of decoupling, the CMB is influenced by the total radiation energy density. Any excess energy that behaves as radiation-like fields may impact the CMB Sendra and Smith (2012). In our case, the primordial GWs carry significant energy, and since GWs evolve as $a^{-4}$ like radiation, frequencies $f>10^{-15}$ Hz can be treated as extra radiation energy that may affect the CMB (see related discussions in Caprini and Figueroa (2018); Clarke et al. (2020b)). This excess radiation can be quantified by the additional relativistic degrees of freedom at decoupling, expressed as $\Delta N_{\text{eff}}$, indicating the extra neutrino species predicted beyond the Standard Model. Using this, we derive the following constraint Caprini and Figueroa (2018); Maiti et al. (2024),
@@
Assuming the present-day photon density parameter as $\Omega_{\gamma}h^{2}\simeq 2.48\times 10^{-5}$, the combined latest Planck-2018 and Baryon Acoustic Oscillation (BAO) data predict $\Delta N_{\text{eff}}\simeq 0.284$ (within a $2\sigma$ range) Aghanim et al. (2020), setting an upper bound on primordial GWs, $\Omega_{\mathrm{gw}}h^{2}<1.67\times 10^{-6}$ Clarke et al. (2020a).
At high frequencies, the GW spectrum behaves as $\Omega_{\mathrm{gw}}h^{2}\propto f^{2n_{\mathrm{B}}}$ for $f<f_{\rm re}$, while for modes well within the horizon $(f\gg f_{\rm re})$, it follows $\Omega_{\mathrm{gw}}h^{2}\propto f^{2n_{\mathrm{B}}-n_{w}}$. Therefore, if $2n_{\mathrm{B}}>n_{w}$, the spectrum has a red tilt for $f>f_{\rm re}$; otherwise, it has a blue tilt. Using this bound, we find that for reheating scenarios with $w_{re}=0$, the maximum allowed magnetic spectral index consistent with the $\Delta N_{\text{eff}}$ bound is $n_{\mathrm{B}}\leq 1.0$. This constraint also restricts the reheating temperature for a fixed magnetic spectral index and equation of state (EoS). For instance, with $w_{re}=0$ and $n_{\mathrm{B}}=1.0$, the reheating temperature should be below $T_{\rm re}<10^{3}$ GeV. For $n_{\mathrm{B}}=0.9$ and $w_{re}=0$, a more relaxed bound is obtained: $T_{\rm re}\leq 10^{7}$ GeV.
For higher EoS, $w_{re}>1/3$, at high frequencies with $n_{\mathrm{B}}>0.3$, the GW spectrum follows $\Omega_{\mathrm{gw}}h^{2}\propto f^{2n_{\mathrm{B}}+|n_{w}|}$. In this regime, secondary GW production during inflation is more substantial compared to late-time production, and since the background dilutes faster than the GW energy density (for $w_{re}>1/3$ scenarios), these frequency modes are more enhanced. Thus, for $w_{re}>1/3$, this bound imposes even stronger constraints on both the magnetic spectral index and reheating dynamics.
@@
As observed, inflationary magnetogenesis models can produce sufficient GW strength to surpass vacuum production in intermediate frequency ranges. For $n_{\mathrm{B}}>0$, with appropriate reheating parameters, we can generate blue-tilted GW at intermediate frequencies. From the $\Delta N_{\text{eff}}$ bound, we found that for $w_{re}>1/3$ reheating scenarios, the magnetic spectral index $n_{\mathrm{B}}$ (or electric spectral index $n_{\mathrm{E}}$) must be tightly constrained to avoid excessive GW production at very high frequencies, where the GW spectrum scales as $\Omega_{\mathrm{gw}}h^{2}\propto f^{2n_{\mathrm{B}}+|n_{w}|}$ for $w_{re}>1/3$. This is illustrated in Fig. (3). Meanwhile, PTA observations indicate that the spectrum should exhibit a strong blue tilt at nano-Hz frequencies, with a GW spectrum scaling as $\Omega_{\mathrm{gw}}h^{2}\propto f^{0.91}$ for NANOGrav 15 yr data Agazie et al. (2023). Here we have found that for $w_{re}<1/3$ reheating scenarios, the GW spectrum produced from the EM field behaves as $\Omega_{\mathrm{gw}}h^{2}\propto f^{2n_{{}_{\mathrm{B}/\mathrm{E}}}-|n_{w}|}$, so to get a strongly blue tilted GW spectrum at the nano-Hz frequency range, the lower EoS $(w_{re}<1/3)$ is more favored compared to $w_{re}>1/3$ scenarios.
For this analysis, we employed the PTARcade code Mitridate et al. (2023) for MCMC analysis to determine the best-fit parameter values, restricting our scenarios to $w_{re}<1/3$ reheating cases and running the PTARcade code. We selected the following priors: $w_{re}\sim\mathcal{U}(0,1/3)$, $n_{\mathrm{B}}\sim\mathcal{U}(0,1.2)$, and $\log_{10}(T_{\rm re})\sim\mathcal{U}(-2,1)$. The predicted parameters and their $1\sigma$ tolerance range, along with the corresponding Bayesian factor, are listed in Tab. (3). The posterior distributions of the predicted parameters are shown in Fig. (5) for two different runs $(R1,R2)$ corresponding to two different priors of $n_{\mathrm{B}}\in(0,1.0)$ and $n_{\mathrm{B}}\in((0,1.5))$ respectively. Our analysis indicates that, to match current data, this model can only produce a GW signal consistent with both amplitude and slope if $n_{\mathrm{B}}\simeq 1.21^{+0.1}_{-0.12}$. Note the $n_{\mathrm{B}}$ is greater than unity.

### v2

@@ III.3.2 Spectral Analysis of Secondary Gravitational Waves (SGWs)
In scenarios with infinite electrical conductivity during reheating, electric fields are rapidly dissipated, implying that any contributions from electric fields to the secondary gravitational wave (SGW) background arise solely from their inflationary production. In contrast, magnetic fields can persist and contribute during both the reheating and radiation-dominated eras. However, as demonstrated, even for reheating scenarios with $w_{re}>1/3$, the magnetic field amplitude remains relatively weak, rendering the induced SGW signal subdominant compared to the primary gravitational waves generated during inflation.
Therefore, for reheating scenarios with $w_{re}<1/3$, the SGW spectrum retains the same spectral behavior as described in Eq. (60). Conversely, for $w_{re}>1/3$, the SGW spectrum is no longer dominated by the magnetic field alone and may receive non-negligible contributions from the residual electric field produced during inflation.
$\Omega_{\mathrm{gw}}^{\mathrm{sec}}(k)h^{2}\simeq$  $\,2.26\times 10^{-26}\left(\frac{\Omega_{R}h^{2}}{4.3\times 10^{-5}}\right)\left(\frac{H_{{}_{\mathrm{I}}}}{10^{-5}M_{\rm P}}\right)^{4}\left(\frac{k_{\mathrm{re}}}{k_{\rm end}}\right)^{2n_{{}_{\mathrm{B}/\mathrm{E}}}}\mathcal{I}_{\rm inf}^{2}\mathcal{C}_{\beta}^{2}(n_{{}_{\mathrm{B}/\mathrm{E}}})\,\mathcal{F}_{\beta}(k)$
$\times\begin{cases}\left(\frac{k}{k_{\mathrm{re}}}\right)^{2n_{\mathrm{E}}},&k_{*}<k<k_{\mathrm{re}},\\ \left(\frac{k}{k_{\mathrm{re}}}\right)^{2n_{\mathrm{E}}+|n_{w}|},&k_{\mathrm{re}}<k<k_{\rm e}.\end{cases}$ (62)
To achieve a successful magnetogenesis scenario without encountering strong coupling or backreaction issues, the zero electrical conductivity limit during reheating is particularly favorable within the coupling range 2.2<n<0-2.2<n<0 (equivalently, $0.4<n_{\mathrm{E}}<4$). In particular, for the subclass with n<2.0n<-2.0 (i.e., $n_{\mathrm{E}}<0$), the electric field produced during inflation is red-tilted. This red tilt enhances the power at large scales and can significantly impact the tensor power spectrum, including at cosmic microwave background (CMB) scales, for both $w_{re}<1/3$ and $w_{re}>1/3$ reheating scenarios.
Most notably, these scenarios can generate a sizable tensor-to-scalar ratio at CMB scales, potentially exceeding the current observational upper bound $r_{0.05}\leq 0.036$. In the following discussion, we explore how this observational constraint can be used to place stringent bounds on both the reheating dynamics and the magnetogenesis model parameter $n$.
Figure 3: In this figure, we plot $\Omega_{\mathrm{gw}}h^{2}$ versus $f$ (in Hz) for $w_{re}=0$ for two different scenarios. In the left panel, we show how the present-day GWs strength depends on the magnetogenesis parameter $n_{\mathrm{E}}(=4-2|n|)$ for a fixed reheating temperature $T_{\rm re}=10^{3}\,\mbox{GeV}$. Here, four different colors indicate four different values os the electric spectral index. Whereas in the right panel, we have shown the dependency of the reheating temperature $T_{\rm re}$, indicated through four different colors with a fixed electric spectral index $n_{\mathrm{E}}=0.6$.
Figure 4: In the figures above, we have plotted $\Omega_{\mathrm{gw}}h^{2}$ as a function of frequency ff(in Hz) for the coupling n<1/2n<-1/2, assuming zero electrical conductivity during the reheating era for $w_{re}=3/5$ (stiff like fluid during reheating). In the left plot, we have shown the dependency of the GW spectrum as a function of the electrical spectral index $n_{\mathrm{E}}$ defined via four different colors for a fixed reheating temperature $T_{\rm re}=10^{6}\,\ GeV$. In the right panel, we have shown how the present-day GW strength and spectral shape are dependent on the reheating temperature, denoted by four different colors for a fixed electric al spectral index $n_{\mathrm{E}}=0.2$.
@@ IV Results and Discussion:
In Fig. 3, we present the present-day gravitational wave (GW) energy spectrum, $\Omega_{\mathrm{gw}}h^{2}(f)$, as a function of the observed frequency $f\,(\mathrm{Hz})$, for two distinct reheating scenarios. In the left panel, we illustrate the impact of the electric spectral index $n_{\mathrm{E}}$ on the GW spectrum, keeping the equation of state fixed at $w_{re}=0$ and the reheating temperature at $T_{\rm re}=10^{3}\,\mathrm{GeV}$. As expected, at very large wavelengths (small $f$), the GW spectrum remains nearly scale-invariant, $\Omega_{\mathrm{gw}}h^{2}(f\ll f_{\rm re})\propto f^{0}$, due to the dominance of primary gravitational waves. This behavior is dictated by the scale-invariant nature of tensor modes produced from quantum fluctuations during inflation. We consider values $n_{\mathrm{E}}=0.4,\,0.5,\,0.6,$ and $0.7$, corresponding to strongly blue-tilted electric spectra. In such cases, most of the electromagnetic energy is concentrated at high frequencies, resulting in significant GW production in that regime. Consequently, in the intermediate frequency range $f_{\mathrm{cmb}}\ll f<f_{\rm re}$, the GW spectrum acquires a blue tilt, while at higher frequencies $f>f_{\rm re}$, the spectrum becomes red-tilted. The spectral slope in this regime can be parametrized as $\Omega_{\mathrm{gw}}h^{2}(f>f_{\mathrm{re}})\propto f^{2n_{\mathrm{E}}-|n_{w}|}$, where $n_{w}=2(1-3w_{re})/(1+3w_{re})$. In the right panel of Fig. 3, we demonstrate how the reheating temperature $T_{\rm re}$ affects the GW spectrum, fixing $n_{\mathrm{E}}=0.6$ and $w_{re}=0$. In both cases, we observe that for modes that enter the horizon before the end of reheating ($f>f_{\rm re}$), the spectrum exhibits a red-tilted behavior, driven by the differential dilution rates of the background energy density and the electric field during reheating.
In Fig. 4, we present the present-day gravitational wave (GW) spectrum $\Omega_{\mathrm{gw}}h^{2}(f)$ as a function of the observable frequency $f$ (in Hz) for reheating scenarios with an equation of state $w_{re}>1/3$. For such scenarios, the spectrum typically exhibits more spectral features compared to the $w_{re}<1/3$ case, owing to the more complex interplay between primary and secondary GW sources.
The left panel illustrates the dependence of the GW spectrum on the electric spectral index $n_{\mathrm{E}}$, fixing $w_{re}=3/5$ and the reheating temperature $T_{\rm re}=10^{6}\,\mathrm{GeV}$. Depending on the value of $n_{\mathrm{E}}$, we observe a variety of spectral structures: in some cases, the spectrum exhibits three distinct breaks, in others two, and occasionally only one. This rich spectral behavior arises from the varying dominance of the different contributions to the GW background.
For instance, considering $n_{\mathrm{E}}=0.15$ with $w_{re}=3/5$ and $T_{\rm re}=10^{6}\,\mathrm{GeV}$, the spectrum near the CMB scale appears scale-invariant due to the dominance of primary GWs (PGWs) over secondary GWs (SGWs). At intermediate frequencies $f<f_{\rm re}$, the spectrum transitions to a blue-tilted behavior $\Omega_{\mathrm{gw}}h^{2}(f)\propto f^{2n_{\mathrm{E}}}$, as SGWs sourced by the blue-tilted electric field become dominant. In the range $f_{\rm re}\ll f\ll f_{\mathrm{end}}$, the spectrum shows a red-tilted behavior due to the suppression of SGWs by reheating dynamics. Finally, at high frequencies $f\gg f_{\rm re}$, the spectrum becomes blue-tilted again, following $\Omega_{\mathrm{gw}}h^{2}(f)\propto f^{|n_{w}|}$, where PGWs regain dominance.
The right panel of Fig. 4 shows the impact of the reheating duration on the present-day GW amplitude. For a fixed $w_{re}=3/5$, increasing the reheating temperature leads to a gradual suppression of the overall GW amplitude. This behavior is consistent with the scaling $(k_{\rm e}/k_{\mathrm{re}})^{2|n_{w}|}$, which parametrizes the enhancement due to electromagnetic field contributions during reheating for $w_{re}>1/3$.
@@
Using the limit $r_{0.05}\leq 0.036$, we find that the coupling parameter $n$ for magnetogenesis should satisfy $2.1\leq n\leq 0$. This bound holds universally, even for instantaneous reheating. For prolonged reheating scenarios with $w_{re}>1/3$, these constraints become more robust, limiting both the reheating dynamics and the magnetogenesis model. Thus, utilizing the upper bound $r_{0.05}\leq 0.036$ at the pivot scale, we derive the following constraint,
$1\geq\frac{182~rA_{s}\mathcal{A}_{1}}{(3-2n_{\mathrm{E}})}\left(\frac{k_{\rm end}}{k_{\mathrm{re}}}\right)^{2|n_{w}|}\left(\frac{k_{*}}{k_{\rm end}}\right)^{2n_{\mathrm{E}}}\mathcal{C}_{\rm E}^{2}(n_{\mathrm{E}}).$ (63)
Using this constraint, we set bounds on the reheating temperature for specific values of the equation of state (EoS) across five discrete values of the magnetic spectral index $n_{\mathrm{E}}=0.01,0.1,0.2,0.3,\text{ and }0.4$, as shown in Table (1).
$w_{re}$ $n_{\mathrm{E}}=0.01$ $n_{\mathrm{E}}=0.1$ $n_{\mathrm{E}}=0.2$ $n_{\mathrm{E}}=0.3$ $n_{\mathrm{E}}=0.4$
@@
0.4 $1.77\times 10^{9}$ - - - -
0.5 $5.65\times 10^{12}$ $1.28\times 10^{7}$ $2.89$ - -
0.6 $8.2\times 10^{13}$ $1.46\times 10^{10}$ $5.41\times 10^{5}$ - -
@@
Table 1: In the table above, we’ve listed the minimum reheating temperature for various equations of states (denoted as $w_{re}$), considering five different values of the electric spectral index (denoted as $n_{\mathrm{E}}$).
In our analysis, we treat the reheating equation-of-state parameter $w_{re}$ as a free parameter, allowing it to take arbitrary values. However, in specific inflationary models—for example, those with power-law potentials of the form $V(\phi)\propto\phi^{2m}$, $w_{re}$ can be uniquely determined based on the dynamics of the scalar field. In this work, we adopt a generalized approach by leaving $w_{re}$ unconstrained. As an illustration, to avoid the overproduction of tensor perturbations at CMB scales, consider the case where the magnetic spectral index is $n_{\mathrm{E}}=0.01$ (nearly scale-invariant). For a reheating equation-of-state $w_{re}=0.4$, the lower bound on the reheating temperature is found to be $T_{\rm re}^{\rm min}\simeq 1.77\times 10^{9}\,\mathrm{GeV}$. If the reheating temperature falls below this value, electromagnetic fields generated via this mechanism can produce excessive tensor fluctuations at large scales, violating current observational constraints on the tensor-to-scalar ratio. We summarize the lower bounds on reheating temperatures for various parameter combinations in Table (1).
@@
During the epoch of decoupling, the CMB is influenced by the total radiation energy density. Any excess energy that behaves as radiation-like fields may impact the CMB Sendra and Smith (2012). In our case, the primordial GWs carry significant energy, and since GWs evolve as $a^{-4}$ like radiation, frequencies $f>10^{-15}$ Hz can be treated as extra radiation energy that may affect the CMB (see related discussions in Caprini and Figueroa (2018); Clarke et al. (2020a)). This excess radiation can be quantified by the additional relativistic degrees of freedom at decoupling, expressed as $\Delta N_{\text{eff}}$, indicating the extra neutrino species predicted beyond the Standard Model. Using this, we derive the following constraint Caprini and Figueroa (2018); Maiti et al. (2024),
@@
Assuming the present-day photon density parameter as $\Omega_{\gamma}h^{2}\simeq 2.48\times 10^{-5}$, the combined latest Planck-2018 and Baryon Acoustic Oscillation (BAO) data predict $\Delta N_{\text{eff}}\simeq 0.284$ (within a $2\sigma$ range) Aghanim et al. (2020), setting an upper bound on primordial GWs, $\Omega_{\mathrm{gw}}h^{2}<1.67\times 10^{-6}$ Clarke et al. (2020b).
At high frequencies, the GW spectrum behaves as $\Omega_{\mathrm{gw}}h^{2}\propto f^{2n_{\mathrm{E}}}$ for $f<f_{\rm re}$, while for modes well within the horizon $(f\gg f_{\rm re})$, it follows $\Omega_{\mathrm{gw}}h^{2}\propto f^{2n_{\mathrm{E}}-n_{w}}$. Therefore, if $2n_{\mathrm{E}}>n_{w}$, the spectrum has a red tilt for $f>f_{\rm re}$; otherwise, it has a blue tilt. Using this bound, we find that for reheating scenarios with $w_{re}=0$, the maximum allowed magnetic spectral index consistent with the $\Delta N_{\text{eff}}$ bound is $n_{\mathrm{E}}\leq 1.0$. This constraint also restricts the reheating temperature for a fixed magnetic spectral index and equation of state (EoS). For instance, with $w_{re}=0$ and $n_{\mathrm{E}}=1.0$, the reheating temperature should be below $T_{\rm re}<10^{3}$ GeV. For $n_{\mathrm{E}}=0.9$ and $w_{re}=0$, a more relaxed bound is obtained, $T_{\rm re}\leq 10^{7}$ GeV.
For higher EoS, $w_{re}>1/3$, at high frequencies with $n_{\mathrm{E}}>0.3$, the GW spectrum follows $\Omega_{\mathrm{gw}}h^{2}\propto f^{2n_{\mathrm{E}}+|n_{w}|}$. In this regime, secondary GW production during inflation is more substantial compared to late-time production, and since the background dilutes faster than the GW energy density (for $w_{re}>1/3$ scenarios), these frequency modes are more enhanced. Thus, for $w_{re}>1/3$, this bound imposes even stronger constraints on both the magnetic spectral index and reheating dynamics.
@@
As observed, inflationary magnetogenesis models can produce sufficient GW strength to surpass vacuum production in intermediate frequency ranges. For $n_{\mathrm{E}}>0$, with appropriate reheating parameters, we can generate blue-tilted GW at intermediate frequencies. From the $\Delta N_{\text{eff}}$ bound, we found that for $w_{re}>1/3$ reheating scenarios, the electrical spectral index $n_{\mathrm{E}}$ must be tightly constrained to avoid excessive GW production at very high frequencies, where the GW spectrum scales as $\Omega_{\mathrm{gw}}h^{2}\propto f^{2n_{\mathrm{E}}+|n_{w}|}$ for $w_{re}>1/3$. This is illustrated in Fig. (4). Meanwhile, PTA observations indicate that the spectrum should exhibit a strong blue tilt at nano-Hz frequencies, with a GW spectrum scaling as $\Omega_{\mathrm{gw}}h^{2}\propto f^{0.91}$ for NANOGrav 15 yr data Agazie et al. (2023). Here we have found that for $w_{re}<1/3$ reheating scenarios, the GW spectrum produced from the EM field behaves as $\Omega_{\mathrm{gw}}h^{2}\propto f^{2n_{\mathrm{E}}}$ for $f<f_{\rm re}$ and $\Omega_{\mathrm{gw}}h^{2}(f)\propto f^{2n_{\mathrm{E}}-|n_{w}|}$ for $f>f_{\rm re}$, so to get a strongly blue tilted GW spectrum at the nano-Hz frequency range, the lower EoS $(w_{re}<1/3)$ is more favored compared to $w_{re}>1/3$ scenarios.
For this analysis, we employed the PTARcade code Mitridate et al. (2023) for Markov Chain Monte Carlo (MCMC) analysis to determine the best-fit parameter values, restricting our scenarios to $w_{re}<1/3$ reheating cases and running the PTARcade code. We selected the following priors: $w_{re}\sim\mathcal{U}(0,1/3)$, $n_{\mathrm{E}}\sim\mathcal{U}(0,1.2)$, and $\log_{10}(T_{\rm re})\sim\mathcal{U}(-2,1)$. The predicted parameters and their $1\sigma$ tolerance range, along with the corresponding Bayesian factor, are listed in Tab. (2). The posterior distributions of the predicted parameters are shown in Fig. (5) for two different runs $(R1,R2)$ corresponding to two different priors of $n_{\mathrm{E}}\in(0,1.0)$ and $n_{\mathrm{E}}\in(0,1.5)$ respectively. Our analysis indicates that, to match current data, this model can only produce a GW signal consistent with both amplitude and slope if $n_{\mathrm{E}}\simeq 1.21^{+0.1}_{-0.12}$. Note the $n_{\mathrm{E}}$ is greater than unity.